Renfro Rentals has issued bonds that have an 11% coupon rate, payable semiannually. The bonds mature in 8 years, have a face value of $1,000, and a yield to maturity of 9.5%. What is the price of the bonds? Round your answer to the nearest cent.
Answer #1

Information provided:

Face value= future value= $1,000

Time= 8 years*2= 16 semi-annual periods

Coupon rate= 11%/2= 5.5%

Coupon payment= 0.055*1,000= $55 per semi-annual period

Yield to maturity= 9.5%/2= 4.75% per semi-annual period

The price of the bond is calculated by computing the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

PMT= 55

I/Y= 4.75

N= 16

Press the CPT key and PV to compute the present value.

The value obtained is 1,082.75

Therefore, the price of the bond is $1,082.75

In case of any query, kindly comment on the solution.